A new factory in a small town has an annual payroll of $6 million. It is expected that 60% of this money will be spent in the town by factory personnel. The people in the town who receive this money are expected to spend 60% of what they receive in the town, and so on. What is the total of all this spending, called the total economic impact of the factory, on the town each year?


1
Expert's answer
2022-01-21T09:16:53-0500

"P=6\cdot(0.6+0.36+0.216+...)=6\cdot\frac{0.6}{1-0.6}=6\cdot\frac{0.6}{0.4}=9" millions.